# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ A.1 Model Formulation
We use a linear Gaussian transition model to describe the weight evolution over time: For each class $k$, the weight vector evolves according to a linear drift parameterized by a class-specific transition matrix ${\bm{A}}_{k}\in{\mathbb{R}}^{D\times D}$. This allows each class to have independent dynamics. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The transition noise covariance matrix is a shared parameter across classes and time points to prevent overfitting and keep parameter size at bay. Eqn. 10 states the Gaussian transition density.
@@
Transition model: $\quad p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}{\mathcal{N}}({\mathbf{w}}_{t,k}|{\bm{A}}_{k}{\mathbf{w}}_{t-1,k},\bm{\Sigma}^{\mathrm{trans}})$ (10)
@@
Eqn. 11 gives the emission model of the observed features ${\mathbf{H}}_{t}$ at time $t$. As in Eqn. 2, the features at a given time $t$ are generated by a mixture distribution with mixing coefficient $\pi_{t,k}$. The emission density of each of the $K$ component is a multivariate normal with the weight vector of class $k$ at time $t$ as mean and $\bm{\Sigma}^{\mathrm{ems}}\in{\mathbb{R}}^{D\times D}$ as class-independent covariance matrix. The resulting model can be seen as a mixture of $K$ Kalman filters. Variants of it has found application in applied statistics (Calabrese & Paninski, 2011).
@@
We use the EM objective of Eqn. 5 to maximize for the model parameters $\phi=\{\{{\bm{A}}_{k},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K},\bm{\Sigma}^{\mathrm{trans}},\bm{\Sigma}^{\mathrm{ems}}\}$. Thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn. 5 can be computed analytically using the well known Kalman filter predict, update and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006).
@@
The closed form computations of the posterior $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$ and smoothing $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:T})$ densities come at a cost as they involve amongst others matrix inversions of dimensionality $D\times D$. This results in considerable computational costs and can lead to numerical instabilities when feature dimension $D$ is large. In addition, the parameter size scales $K\times D^{2}$ risking overfitting and consuming substantial memory. These are limitations of the Gaussian formulation making it costly for high-dimensional feature spaces and impractical in low resource environments requiring instant predictions.
Appendix B Details on STAD-vMF
B.1 Inference
@@
Using the von Mises-Fisher distribution as hyperspherical transition and emission model, the log of the complete-data likelihood in Eqn. 3 becomes
$\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=$ $\sum_{k}^{K}\log p({\mathbf{w}}_{1,k})$ (12)
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\log p({\mathbf{c}}_{t,n})+\sum_{k=1}^{K}{\textnormal{c}}_{t,n,k}\log p({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$ (13)
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log p({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$ (14)
$=$ $\sum_{k}^{K}\log C_{D}(\kappa_{0,k})+\kappa_{0,k}\bm{\mu}_{0,k}^{T}{\mathbf{w}}_{1,k}$ (15)
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\sum_{k=1}^{K}c_{n,t,k}\big{(}\log\pi_{t,k}+\log C_{D}(\kappa^{\mathrm{ems}})+\kappa^{\mathrm{ems}}{\mathbf{w}}_{t,k}^{T}{\mathbf{h}}_{t,n}\big{)}$ (16)
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log C_{D}(\kappa^{\mathrm{trans}})+\kappa^{\mathrm{trans}}{\mathbf{w}}_{t-1,k}^{T}{\mathbf{w}}_{t,k}$ (17)
where $\kappa_{0,k}$ and $\bm{\mu}_{0,k}$ denote the parameters of the first time step. In practise, we set $\bm{\mu}_{0,k}$ to the source weights and $\kappa_{0,k}=100$ (see App. C).
@@
As described in Sec. 3.3, we approximate the posterior $p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$ with a variational distribution $q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})$ assuming the factorised form
$q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=\prod_{t=1}^{T}\prod_{k=1}^{K}q({\mathbf{w}}_{t,k})\prod_{n=1}^{N_{t}}q({\mathbf{c}}_{n,t}),$ (18)
where we parameterise $q({\mathbf{w}}_{t,k})$ and $q({\mathbf{c}}_{n,t})$ with
$q({\mathbf{w}}_{t,k})={\text{{vMF}}}(\ \cdot\ ;\bm{\rho_{t,k}},\gamma_{t,k})\quad q({\mathbf{c}}_{n,t})=\text{{Cat}}(\ \cdot\ ;\bm{\lambda_{n,t}})\quad\forall t,n,k.$ (19)
@@
Taking the expectation of the complete-data log likelihood (Eqn. 12) with respect to the variational distribution (Eqn. 18) gives
$\mathbb{E}_{q}[\log p({\mathbf{H}}_{1:T},$ ${\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})]=\sum_{k}^{K}\log C_{D}(\kappa_{0,k})+\kappa_{0,k}\bm{\mu}_{0,k}^{T}\mathbb{E}_{q}[{\mathbf{w}}_{1,k}]$ (21)
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\sum_{k=1}^{K}\mathbb{E}_{q}[c_{n,t,k}]\big{(}\log\pi_{t,k}+\log C_{D}(\kappa^{\mathrm{ems}})+\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{T}{\mathbf{h}}_{t,n}\big{)}$ (22)
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log C_{D}(\kappa^{\mathrm{trans}})+\kappa^{\mathrm{trans}}\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]^{T}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]$ (23)
@@
$\lambda_{n,t,k}=\frac{\beta_{n,t,k}}{\sum_{j=1}^{K}\beta_{n,t,j}}\quad\text{with}\quad\beta_{n,t,k}=\pi_{t,k}C_{D}(\kappa^{\mathrm{ems}})\exp(\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{T}{\mathbf{h}}_{n,t})$ (24)
$\bm{\rho}_{t,k}=\frac{\kappa^{\mathrm{trans}}\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]+\kappa^{\mathrm{ems}}\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[c_{n,t,k}]{\mathbf{h}}_{n,t}+\kappa^{\mathrm{trans}}\mathbb{E}_{q}[{\mathbf{w}}_{t+1,k}]}{\gamma_{t,k}}$ (25)
$\gamma_{t,k}=||\bm{\rho}_{t,k}||$ (26)
@@
$\mathbb{E}[c_{n,t,k}]$ $=\lambda_{n,t,k}$ (27)
$\mathbb{E}[{\mathbf{w}}_{t,k}]$ $=A_{D}(\gamma_{t,k})\bm{\rho_{t,k}},$ (28)
@@
Maximizing objective (Eqn. 20) with respect to the model parameters $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ gives
$\hat{\kappa}^{\mathrm{trans}}=$ $\frac{\bar{r}^{\mathrm{trans}}D-(\bar{r}^{\mathrm{trans}})^{3}}{1-(\bar{r}^{\mathrm{trans}})^{2}}\quad\text{with}\quad\bar{r}^{\mathrm{trans}}=\left\|\frac{\sum_{t=2}^{T}\sum_{k=1}^{K}\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]^{T}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]}{(T-1)\times K}\right\|$ (29)
$\hat{\kappa}^{\mathrm{ems}}=$ $\frac{\bar{r}^{\mathrm{ems}}D-(\bar{r}^{\mathrm{ems}})^{3}}{1-(\bar{r}^{\mathrm{ems}})^{2}}\quad\text{with}\quad\bar{r}^{\mathrm{ems}}=\left\|\frac{\sum_{t=2}^{T}\sum_{k=1}^{K}\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[{\mathbf{c}}_{n,t,k}]\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{T}{\mathbf{h}}_{n,t}}{\sum_{t=1}^{T}N_{t}}\right\|$ (30)
$\pi_{t,k}=$ $\frac{\sum_{n=1}^{N_{t}}\mathbb{E}[c_{n,t,k}]}{N_{t}}$ (31)

### v3

@@ C.1 Details on STAD-Gaussian
We use a linear Gaussian transition model to describe the weight evolution over time: For each class $k$, the weight vector evolves according to a linear drift parameterized by a class-specific transition matrix $\mathbf{A}_{k}\in{\mathbb{R}}^{D\times D}$. This allows each class to have independent dynamics. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The transition noise covariance matrix is a shared parameter across classes and time points to prevent overfitting and keep the parameter size manageable. Eqn.˜11 states the Gaussian transition density.
@@
Transition model: $\quad p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1})=\prod_{k=1}^{K}\mathcal{N}({\mathbf{w}}_{t,k}|\mathbf{A}_{k}{\mathbf{w}}_{t-1,k},\bm{\Sigma}^{\mathrm{trans}})$ (11)
@@
Eqn.˜12 gives the emission model of the observed features ${\mathbf{H}}_{t}$ at time $t$. As in Eqn.˜2, the features at a given time $t$ are generated by a mixture distribution with mixing coefficient $\pi_{t,k}$. The emission density of each of the $K$ components is a multivariate normal with the weight vector of class $k$ at time $t$ as mean and $\bm{\Sigma}^{\mathrm{ems}}\in{\mathbb{R}}^{D\times D}$ as class-independent covariance matrix. The resulting model can be seen as a mixture of $K$ Kalman filters. Variants of it have found application in applied statistics (Calabrese & Paninski, 2011).
@@
We use the EM objective of Eqn.˜4 to maximize for the model parameters $\phi=\{\{\mathbf{A}_{k},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K},\bm{\Sigma}^{\mathrm{trans}},\bm{\Sigma}^{\mathrm{ems}}\}$. Thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn.˜4 (E-step) can be computed analytically using the well-known Kalman filter predict, update, and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006) as outlined next.
E-step
The posterior responsibilities of each component are given by the Gaussian mixture posterior:
$\mathbb{E}[c_{t,n,k}]=\frac{\pi_{t,k}|\bm{\Sigma}^{\mathrm{ems}}|^{-1/2}\exp\!\left\{-\frac{1}{2}({\mathbf{h}}_{t,n}-{\mathbf{w}}_{t,k})^{\top}(\bm{\Sigma}^{\mathrm{ems}})^{-1}({\mathbf{h}}_{t,n}-{\mathbf{w}}_{t,k})\right\}}{\sum_{j=1}^{K}\pi_{t,j}|\bm{\Sigma}^{\mathrm{ems}}|^{-1/2}\exp\!\left\{-\frac{1}{2}({\mathbf{h}}_{t,n}-{\mathbf{w}}_{t,j})^{\top}(\bm{\Sigma}^{\mathrm{ems}})^{-1}({\mathbf{h}}_{t,n}-{\mathbf{w}}_{t,j})\right\}},$ (13)
which we summarize as $\mathbb{E}[c_{t,n,k}]=\text{computeAssignments}({\mathbf{h}}_{t,n},{\mathbf{w}}_{t,k},\pi_{t,k},\bm{\Sigma}^{\mathrm{ems}})$. Given the previous posterior $(\bm{\mu}_{t-1,k}^{+},\bm{\Sigma}_{t-1,k}^{+})$, the prior distribution at time $t$ follows the standard Kalman filter prediction:
$\bm{\mu}_{t,k}^{-}=\mathbf{A}_{k}\bm{\mu}_{t-1,k}^{+},\quad\bm{\Sigma}_{t,k}^{-}=\bm{\Sigma}^{\mathrm{trans}}+\mathbf{A}_{k}\bm{\Sigma}_{t-1,k}^{+}\mathbf{A}_{k}^{\top}.$ (14)
We denote this step as $(\bm{\mu}_{t,k}^{-},\bm{\Sigma}_{t,k}^{-})=\text{predict}(\mathbf{A}_{k},\bm{\Sigma}^{\mathrm{trans}},\bm{\mu}_{t-1,k}^{+},\bm{\Sigma}_{t-1,k}^{+})$. Conditioning the predicted state on the current observation ${\mathbf{h}}_{t,n}$ yields:
$\bm{\mu}_{t,n,k}^{+}$ $=\bm{\Sigma}_{t,n,k}^{+}\!\left[(\bm{\Sigma}_{t,k}^{-})^{-1}\bm{\mu}_{t,k}^{-}+\mathbb{E}[c_{t,n,k}]\,(\bm{\Sigma}^{\mathrm{ems}})^{-1}{\mathbf{h}}_{t,n}\right],$ (15)
$\bm{\Sigma}_{t,n,k}^{+}$ $=\left[(\bm{\Sigma}_{t,k}^{-})^{-1}+\mathbb{E}[c_{t,n,k}]\,(\bm{\Sigma}^{\mathrm{ems}})^{-1}\right]^{-1},$ (16)
that is, $(\bm{\mu}_{t,n,k}^{+},\bm{\Sigma}_{t,n,k}^{+})=\text{update}(\bm{\mu}_{t,k}^{-},\bm{\Sigma}_{t,k}^{-},{\mathbf{h}}_{t,n},\mathbb{E}[c_{t,n,k}])$. This corresponds to a Kalman filter trajectory for each observation and class. In order to aggregate the observation-wise posterior to one single posterior over each class, we chose a mixture distribution with weights $\alpha_{t,n,k}=\frac{\mathbb{E}[c_{t,n,k}]}{\sum_{i=1}^{N_{t}}\mathbb{E}[c_{t,i,k}]}$ which gives:
$\bm{\mu}_{t,k}^{+}=\sum_{n=1}^{N_{t}}\alpha_{t,n,k}\,\bm{\mu}_{t,n,k}^{+},\quad\bm{\Sigma}_{t,k}^{+}=\sum_{n=1}^{N_{t}}\alpha_{t,n,k}\,\bm{\Sigma}_{t,n,k}^{+}+\sum_{n=1}^{N_{t}}\alpha_{t,n,k}(\bm{\mu}_{t,n,k}^{+}-\bm{\mu}_{t,k}^{+})(\bm{\mu}_{t,n,k}^{+}-\bm{\mu}_{t,k}^{+})^{\top},$ (17)
or compactly $(\bm{\mu}_{t,k}^{+},\bm{\Sigma}_{t,k}^{+})=\text{mixture}(\{\bm{\mu}_{t,n,k}^{+},\bm{\Sigma}_{t,n,k}^{+},\alpha_{t,n,k}\}_{n=1}^{N_{t}})$. Smoothing backward in time gives:
$\bm{\mu}_{t,k}$ $=\bm{\mu}_{t,k}^{+}+\mathbf{J}_{t,k}\!\left(\bm{\mu}_{t+1,k}-\mathbf{A}_{k}\bm{\mu}_{t,k}^{+}\right),\quad\bm{\Sigma}_{t,k}=\bm{\Sigma}_{t,k}^{+}+\mathbf{J}_{t,k}\!\left(\bm{\Sigma}_{t+1,k}-\bm{\Sigma}_{t,k}^{-}\right)\mathbf{J}_{t,k}^{\top},$ (18)
$\mathbf{J}_{t,k}$ $=\bm{\Sigma}_{t,k}^{+}\mathbf{A}_{k}^{\top}(\bm{\Sigma}_{t,k}^{-})^{-1},$ (19)
i.e., $(\bm{\mu}_{t,k},\bm{\Sigma}_{t,k})=\text{smooth}(\bm{\mu}_{t,k}^{+},\bm{\Sigma}_{t,k}^{+},\bm{\mu}_{t+1,k},\bm{\Sigma}_{t+1,k},\bm{\Sigma}_{t,k}^{-},\mathbf{A}_{k})$. Finally, from the smoothed posteriors we compute the necessary expectations required in the M-step:
$\mathbb{E}[{\mathbf{w}}_{t,k}]=\bm{\mu}_{t,k},\quad\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}]=\bm{\Sigma}_{t,k}\mathbf{J}_{t-1,k}^{\top}+\bm{\mu}_{t,k}\bm{\mu}_{t-1,k}^{\top},\quad\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}]=\bm{\Sigma}_{t,k}+\bm{\mu}_{t,k}\bm{\mu}_{t,k}^{\top},$ (20)
in short, $\{\mathbb{E}[{\mathbf{w}}_{t,k}],\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}],\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}]\}=\text{computeMarginals}(\bm{\mu}_{t,k},\bm{\Sigma}_{t,k},\bm{\mu}_{t-1,k},\mathbf{J}_{t-1,k})$.
M-step
Given the responsibilities $\mathbb{E}[c_{t,n,k}]$ and the smoothed posteriors of the prototypes $(\bm{\mu}_{t,k},\bm{\Sigma}_{t,k})$, the parameters are re-estimated by maximizing the expected complete-data log-likelihood.
The posterior mixture weights are given by the average responsibilities:
$\pi_{t,k}=\frac{1}{N_{t}}\sum_{n=1}^{N_{t}}\mathbb{E}[c_{t,n,k}],$ (21)
i.e., $\pi_{t,k}=\text{updatePi}(\{\mathbb{E}[c_{t,n,k}]\}_{n=1}^{N_{t}})$. Maximizing the likelihood of the linear-Gaussian state evolution yields the least-squares estimate of the transition matrices:
$\mathbf{A}_{k}=\left(\sum_{t=2}^{T}\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}]\right)\left(\sum_{t=2}^{T}\mathbb{E}[{\mathbf{w}}_{t-1,k}{\mathbf{w}}_{t-1,k}^{\top}]\right)^{-1},$ (22)
or equivalently $\mathbf{A}_{k}=\text{updateA}(\{\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}],\mathbb{E}[{\mathbf{w}}_{t-1,k}{\mathbf{w}}_{t-1,k}^{\top}]\}_{t=2}^{T})$. The transition noise covariance is obtained by the maximum-likelihood estimator of the process noise:
$\bm{\Sigma}^{\mathrm{trans}}=\frac{1}{(T-1)K}\sum_{k=1}^{K}\sum_{t=2}^{T}\Big(\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}]-\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}]\mathbf{A}_{k}^{\top}-\mathbf{A}_{k}\mathbb{E}[{\mathbf{w}}_{t-1,k}{\mathbf{w}}_{t,k}^{\top}]+\mathbf{A}_{k}\mathbb{E}[{\mathbf{w}}_{t-1,k}{\mathbf{w}}_{t-1,k}^{\top}]\mathbf{A}_{k}^{\top}\Big),$ (23)
in short, $\bm{\Sigma}^{\mathrm{trans}}=\text{updateSigmaTrans}(\{\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}],\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t-1,k}^{\top}],\mathbb{E}[{\mathbf{w}}_{t-1,k}{\mathbf{w}}_{t-1,k}^{\top}],\mathbf{A}_{k}\}_{t=2}^{T})$. Finally, maximizing the likelihood of the Gaussian emission model yields the responsibility-weighted residual covariance:
$\bm{\Sigma}^{\mathrm{ems}}=\frac{\sum_{k=1}^{K}\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\mathbb{E}[c_{t,n,k}]\big({\mathbf{h}}_{t,n}{\mathbf{h}}_{t,n}^{\top}-\mathbb{E}[{\mathbf{w}}_{t,k}]{\mathbf{h}}_{t,n}^{\top}-{\mathbf{h}}_{t,n}\mathbb{E}[{\mathbf{w}}_{t,k}]^{\top}+\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}]\big)}{\sum_{k=1}^{K}\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\mathbb{E}[c_{t,n,k}]},$ (24)
or compactly $\bm{\Sigma}^{\mathrm{ems}}=\text{updateSigmaEms}(\{{\mathbf{h}}_{t,n},\mathbb{E}[c_{t,n,k}],\mathbb{E}[{\mathbf{w}}_{t,k}],\mathbb{E}[{\mathbf{w}}_{t,k}{\mathbf{w}}_{t,k}^{\top}]\}_{t,n,k})$.
@@
The closed-form computations of the posterior $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$ and smoothing $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:T})$ densities come at a cost as they involve, amongst others, matrix inversions of dimensionality $D\times D$. This results in considerable computational costs and can lead to numerical instabilities when feature dimension $D$ is large. In addition, the parameter size scales as $K\times D^{2}$, risking overfitting and consuming substantial memory. These are limitations of the Gaussian formulation, making it costly for high-dimensional feature spaces and impractical in low-resource environments requiring instant predictions.
Algorithm 2 EM for STAD-Gauss
1:Input: $\{{\mathbf{W}}_{\tau},{\mathbf{C}}_{\tau},\bm{\pi}_{\tau},{\mathbf{H}}_{\tau}\}_{\tau\in S_{t}}$, $\bm{\Sigma}^{\mathrm{trans}},\mathbf{A}_{k},\bm{\Sigma}^{\mathrm{ems}}$
2:for $\tau\in S_{t}$ do
3:  E-step
4:  $\mathbb{E}[c_{\tau,n,k}]=\text{computeAssignments}({\mathbf{h}}_{\tau,n},{\mathbf{w}}_{\tau,k},\pi_{\tau,k},\bm{\Sigma}^{\mathrm{ems}})$
5:  $\bm{\mu}_{\tau,k}^{-},\bm{\Sigma}_{\tau,k}^{-}=\text{predict}(\mathbf{A}_{k},\bm{\Sigma}^{\mathrm{trans}},\bm{\mu}_{\tau-1,k}^{+},\bm{\Sigma}_{\tau-1,k}^{+})$
6:  $\bm{\mu}_{\tau,n,k}^{+},\bm{\Sigma}_{\tau,n,k}^{+}=\text{update}(\bm{\mu}_{\tau,k}^{-},\bm{\Sigma}_{\tau,k}^{-},{\mathbf{h}}_{\tau,n},\mathbb{E}[c_{\tau,n,k}])$
7:  $\bm{\mu}_{\tau,k}^{+},\bm{\Sigma}_{\tau,k}^{+}=\text{mixture}(\{\bm{\mu}_{\tau,n,k}^{+},\bm{\Sigma}_{\tau,n,k}^{+},\mathbb{E}[c_{\tau,n,k}]\}_{n=1}^{N_{\tau}})$
8:  $\bm{\mu}_{\tau,k},\bm{\Sigma}_{\tau,k}=\text{smooth}(\bm{\mu}_{\tau,k}^{+},\bm{\Sigma}_{\tau,k}^{+},\bm{\mu}_{\tau+1,k},\bm{\Sigma}_{\tau+1,k},\mathbf{A}_{k})$
9:  $\mathbb{E}[{\mathbf{w}}_{\tau,k}],\mathbb{E}[{\mathbf{w}}_{\tau,k}{\mathbf{w}}_{\tau-1,k}^{\top}],\mathbb{E}[{\mathbf{w}}_{\tau,k}{\mathbf{w}}_{\tau,k}^{\top}]=\text{computeMarginals}(\bm{\mu}_{\tau,k},\bm{\Sigma}_{\tau,k},\bm{\mu}_{\tau-1,k},\mathbf{J}_{\tau-1,k})$
10:  M-step
11:  $\pi_{\tau,k}=\text{updatePi}(\{\mathbb{E}[c_{\tau,n,k}]\}_{n=1}^{N_{\tau}})$
12:end for
[…]
16:return ${\mathbf{W}}_{t},{\mathbf{C}}_{t},\bm{\pi}_{t},\bm{\Sigma}^{\mathrm{trans}},\bm{\Sigma}^{\mathrm{ems}}$
C.2 Details on STAD-vMF
@@
Using the von Mises–Fisher distribution as the hyperspherical transition and emission model, the log of the complete-data likelihood in Eqn.˜3 becomes
$\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=$ $\sum_{k=1}^{K}\log p({\mathbf{w}}_{1,k})$ (25)
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\log p({\mathbf{c}}_{t,n})+\sum_{k=1}^{K}c_{t,n,k}\log p({\mathbf{h}}_{t,n}\mid{\mathbf{w}}_{t,k},\kappa^{\mathrm{ems}})$ (26)
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log p({\mathbf{w}}_{t,k}\mid{\mathbf{w}}_{t-1,k},\kappa^{\mathrm{trans}})$ (27)
$=$ $\sum_{k=1}^{K}\log C_{D}(\kappa_{0,k})+\kappa_{0,k}\bm{\mu}_{0,k}^{\top}{\mathbf{w}}_{1,k}$ (28)
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\sum_{k=1}^{K}c_{t,n,k}\big(\log\pi_{t,k}+\log C_{D}(\kappa^{\mathrm{ems}})+\kappa^{\mathrm{ems}}{\mathbf{w}}_{t,k}^{\top}{\mathbf{h}}_{t,n}\big)$ (29)
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log C_{D}(\kappa^{\mathrm{trans}})+\kappa^{\mathrm{trans}}{\mathbf{w}}_{t-1,k}^{\top}{\mathbf{w}}_{t,k}$ (30)
where $\kappa_{0,k}$ and $\bm{\mu}_{0,k}$ denote the prior parameters for $t=1$. In practice, we set $\bm{\mu}_{0,k}$ to the source weights and $\kappa_{0,k}=100$ (see App.˜D).
@@
As described in Sec.˜3.3, we approximate the posterior $p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}\mid{\mathbf{H}}_{1:T})$ with a variational distribution $q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})$ assuming the factorized form
$q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=\prod_{t=1}^{T}\prod_{k=1}^{K}q({\mathbf{w}}_{t,k})\prod_{n=1}^{N_{t}}q({\mathbf{c}}_{t,n}),$ (31)
where we parameterize $q({\mathbf{w}}_{t,k})$ and $q({\mathbf{c}}_{t,n})$ with
$q({\mathbf{w}}_{t,k})={\text{{vMF}}}(\cdot;\bm{\rho}_{t,k},\gamma_{t,k}),\quad q({\mathbf{c}}_{t,n})=\text{{Cat}}(\cdot;\bm{\lambda}_{t,n}),\quad\forall t,n,k.$ (32)
@@
Taking the expectation of the complete-data log likelihood (Eqn.˜25) with respect to the variational distribution (Eqn.˜31) gives
$\mathbb{E}_{q}$ $[\log p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})]=\sum_{k=1}^{K}\log C_{D}(\kappa_{0,k})+\kappa_{0,k}\bm{\mu}_{0,k}^{\top}\mathbb{E}_{q}[{\mathbf{w}}_{1,k}]$
$+\sum_{t=1}^{T}\sum_{n=1}^{N_{t}}\sum_{k=1}^{K}\mathbb{E}_{q}[c_{t,n,k}]\big(\log\pi_{t,k}+\log C_{D}(\kappa^{\mathrm{ems}})+\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{\top}{\mathbf{h}}_{t,n}\big)$
$+\sum_{t=2}^{T}\sum_{k=1}^{K}\log C_{D}(\kappa^{\mathrm{trans}})+\kappa^{\mathrm{trans}}\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]^{\top}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]$ (34)
@@
$\lambda_{t,n,k}$ $=\frac{\pi_{t,k}C_{D}(\kappa^{\mathrm{ems}})\exp\big\{\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{\top}{\mathbf{h}}_{t,n}\big\}}{\sum_{j=1}^{K}\pi_{t,j}C_{D}(\kappa^{\mathrm{ems}})\exp\big\{\kappa^{\mathrm{ems}}\mathbb{E}_{q}[{\mathbf{w}}_{t,j}]^{\top}{\mathbf{h}}_{t,n}\big\}},$ (35)
[…]
$\beta_{t,k}$ $=\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{t>1\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]+\kappa^{\mathrm{ems}}\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[c_{t,n,k}]{\mathbf{h}}_{t,n}+\kappa^{\mathrm{trans}}\,\mathbb{I}_{\{t<T\}}\,\mathbb{E}_{q}[{\mathbf{w}}_{t+1,k}],$ (37)
@@
$\mathbb{E}_{q}[c_{t,n,k}]$ $=\lambda_{t,n,k}$ (38)
$\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]$ $=A_{D}(\gamma_{t,k})\,\bm{\rho}_{t,k},$ (39)
@@
Maximizing objective (Eqn.˜33) with respect to the model parameters $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\bm{\pi}_{1:T}\}$ gives
$\hat{\kappa}^{\mathrm{trans}}=$ $\frac{\bar{r}^{\mathrm{trans}}D-(\bar{r}^{\mathrm{trans}})^{3}}{1-(\bar{r}^{\mathrm{trans}})^{2}},\quad\text{with}\quad\bar{r}^{\mathrm{trans}}=\left\lVert\frac{\sum_{t=2}^{T}\sum_{k=1}^{K}\mathbb{E}_{q}[{\mathbf{w}}_{t-1,k}]^{\top}\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]}{(T-1)\times K}\right\rVert$ (40)
$\hat{\kappa}^{\mathrm{ems}}=$ $\frac{\bar{r}^{\mathrm{ems}}D-(\bar{r}^{\mathrm{ems}})^{3}}{1-(\bar{r}^{\mathrm{ems}})^{2}},\quad\text{with}\quad\bar{r}^{\mathrm{ems}}=\left\lVert\frac{\sum_{t=1}^{T}\sum_{k=1}^{K}\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[c_{t,n,k}]\,\mathbb{E}_{q}[{\mathbf{w}}_{t,k}]^{\top}{\mathbf{h}}_{t,n}}{\sum_{t=1}^{T}N_{t}}\right\rVert$ (41)
$\pi_{t,k}$ $=\frac{\sum_{n=1}^{N_{t}}\mathbb{E}_{q}[c_{t,n,k}]}{N_{t}}$ (42)